# Regularized Langevin Dynamics for Combinatorial Optimization

### v1

@@ A.3 Postprocessing
We postprocess the best solutions (with the lowest energy) on MIS and MCl problems to satisfy the constraint. We find our method almost surely yields a valid solution, so we just adopt the simplest way to greedily decode it. On MIS, we sort all nodes according to its value (00 or $1$) and initialize the candidate set with all nodes. Each time we select a node, we remove itself and all its neighbors from the candidate set and repeat the process until no candidates are available. On MCl, we perform a similar process by selecting a node and removing itself and all the nodes not in its neighborhood set. on MCut, we simply return the best solution since the problem is not constrained.
Appendix B Comparison under More Running Steps
Here we compare iSCO (Sun et al., 2023) and RLSA by running both methods with 10×10$\times$10 × more steps in Table 6 and 7. Although iSCO has achieved a very close performance to RLSA in some small-scale datasets, RLSA still maintains a clear advantage on large-scale datasets, with significantly less running time.
In particular, we note that RLSA has already achieved comparable performance to the exact solvers over multiple benchmarks in Table 1 and 2, especially on the large-scale problems. This has demonstrated the strong power of RLSA in CO.
Table 6: Comparative results between iSCO and RLSA with $10$ times more steps on MIS. The best one is bolded.
@@ Appendix B Comparison under More Running Steps
Table 7: Comparative results between iSCO and RLSA with $10$ times more steps on MCl and MCut. The best one is bolded.
@@ Appendix C Example Code for RLSA
The following Python code outlines our implementation of RLSA. The energy function corresponds to the formulas in Section 3.2 and the input parameters are in summarized in Section A. In our experiments, the time measurement corresponds to the running time of the entire RLSA function below.

### v2

@@ A.3 Postprocessing
We postprocess the lowest‐energy solutions to enforce feasibility, though in practice our methods almost always produce valid solutions, so we use a simple greedy decoder.
Specifically, we calculate the gradient of the current solution $\mathbf{x}$: if $\max\Delta_{i}<0$, then the solution has already reached a local optimum and the feasibility is guaranteed; otherwise, we flip the coordinate corresponding to $\arg\max\Delta_{i}$. This process is iterated until the convergence.
Appendix B Comparison under Longer Runtime
To assess whether RLSA retains its advantage when given longer runtime, we run both RLSA and iSCO for ten times more iterations than those specified in Table 4. The results in Tables 6 and 7 show that, although iSCO matches RLSA on some small‐scale instances, RLSA still outperforms on large‐scale datasets while requiring substantially less computation time.
Furthermore, the results demonstrate that RLSA achieves performance comparable to exact solvers across multiple benchmarks—especially on large‐scale problem instances—underscoring its effectiveness in CO.
Table 6: Comparative results between iSCO and RLSA with ten times more steps on MIS. The best one is bolded.
@@ Appendix B Comparison under Longer Runtime
Table 7: Comparative results between iSCO and RLSA with ten times more steps on MCl and MCut. The best one is bolded.
@@ Appendix C Example Code for RLSA
The following Python code outlines our implementation of RLSA. The energy function corresponds to the formulas in Section 4.2 and the input parameters are summarized in Section A. In our experiments, the time measurement corresponds to the runtime of the entire RLSA function below.